Paul Coddington Northeast Parallel Architectures Center 111 College Place, 3-228 CST Syracuse, NY 13244

Phone: (315) 443-4883 FAX: (315) 443-1973 Email: paulc@npac.syr.edu

December 2, 1994

Patty Campbell Sailfish Systems 150 Broadway Suite 710 New York, NY 10038

Dear Patty,

Here are the receipts for our travel for the two meetings on September 9 and November 21 of this year. The total cost was $468. Please make the check out to Paul Coddington, since I purchased the tickets.

Thanks,

Paul Coddington